OpenStudy (anonymous):

We measured a 3-level quantum system: \[\frac{1} {\sqrt{3}} |0>+\frac{1} {\sqrt{3}} |1>+\frac{1} {\sqrt{3}} |2>\] with respect to the observable: \[\left[\begin{matrix}1 & 0 & 0\\ 0 & 4 & 0 \\ 0 & 0 & 4\end{matrix}\right]\] and saw a 4. What is the state of the new system?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

So I know we would get the vectors: \[\frac{1}{\sqrt{3}} |1>+\frac{1}{\sqrt{3}} |2>\] but now I'm suppose to renormalize it into a proper quantum state and somehow end up with: \[\frac{1}{\sqrt{2}} |1>+\frac{1}{\sqrt{2}} |2>\] How do I renormalize it?

OpenStudy (anonymous):

Alright, I believe it's because the new state is |+> which translates to the 1/sqrt(2) part. If anyone knows for sure, please let me know.

OpenStudy (unklerhaukus):

the vector \(X\) is normalised \(\hat X\) , when its of unit length \[X=\frac{1}{\sqrt{3}} |1\rangle +\frac{1}{\sqrt{2}} |2\rangle\qquad\qquad|X|=\sqrt{\left(\frac1{\sqrt 3}\right)^2+\left( \frac{2}{\sqrt 2}\right)^2}=\sqrt{\frac{7}{3}}\] \[\hat X=\frac X{|X|}\] check the vector is normalized \[\langle \hat X|\hat X\rangle =1\]